Please show your work with the formulas written out. Please answer as if you can not use a calculator, or only use a four function calculator (because that is how I have to learn it). Do not just put down the calculator keystrokes I need to see every step and number to learn how to do it.

3. What is the future value after 22 years of year-end deposits $250 earning 3%

4. What is the future value in the preceding problem if deposits are year-beginning?
Answer #1

a.

Future Value of Annuity = P/r[(1 + r)n - 1]

Future Value of Annuity = 250/0.03[(1.03)22 - 1]

Future Value = $7,634.20

b.

Future Value of Annuity Due = (1 + r)P/r[(1 + r)n - 1]

Future Value = (1.03)250/0.03[(1.03)22 - 1]

Future Value = $7,863.22